# Towards Imperceptible Backdoor Attack in Self-supervised Learning

### v1

@@ 5.2 Effectiveness Evaluation
Effective Attack.Table 1 indicates our method consistently achieves high ASRs (most above 98%) across various datasets. For instance, with STL10 as the pre-training dataset and CIFAR10 as the downstream dataset, our method reaches an ASR of 99.58%, indicating a highly successful rate of triggering the desired misclassification. This is significantly higher than the ASRs of other benchmark methods like WaNet [51] and CTRL [37] in the same context.
Accuracy Preservation.Table 1 proves that IMPERATIVE preserves the accuracy of the downstream classifiers. The differences between BA and CA are within 2% across the different combinations of pre-training and downstream datasets. This suggests that the presence of the backdoor introduced by our method does not compromise the classifier’s ability to correctly label clean images. For instance, when STL10 is used for pre-training and GTSRB is the downstream dataset, the BA is 76.12% compared to the CA of 75.82%, showing a negligible impact on the classifier’s performance.
@@
5.3 Stealthiness Evaluation
As shown in Table 3, our method exhibits strong stealthiness advantages with SSIM values consistently close to 1 and most above 0.95, indicating minimal structural changes to the images. PSNR values are notably high, exceeding 40 for some cases, which implies that the noise introduced is hardly detectable. Furthermore, the LPIPS values are exceptionally low, ranging from 0.0001 to 0.02, demonstrating that the perceptual difference between the original and perturbed images is almost negligible. These figures collectively suggest that our method introduces changes that are virtually imperceptible to both human eyes and standard image quality metrics.
@@ 5.4 Robustness Evaluation
To assess the resilience of our method against current backdoor defenses, we deploy various state-of-the-art backdoor defense strategies, i.e., DECREE [15], Beatrix [48], Neural Cleanse [83], STRIP [16], Grad-CAM [64] and measure how effectively our method withstands these defenses. Additionally, to further test the robustness of our method, we evaluate its endurance against the following commonly studied noises, i.e., JPEG compression [10, 12, 20, 14], Poisson noise [1, 91], and Salt&Pepper noise [38, 1, 91]. We show the defense results of STRIP, Grad-CAM, and noises in Appendix C.1.
@@
DECREE. [15]  We initially assess whether our method can elude detection by DECREE [15], a method for identifying Trojan attacks in pre-trained encoders. It judges an encoder to be compromised if the proportion of the reversed trigger’s $\mathcal{L}^{1}$ norm relative to the greatest $\mathcal{L}^{1}$ norm within the input space of the model falls below a threshold, i.e., 0.1 $\mathcal{P}\mathcal{L}^{1}$ Norm [15]. As shown in Table 3, $\mathcal{P}\mathcal{L}^{1}$-Norm of each pre-trained and downstream dataset pair are above a default threshold 0.1, and thus this approach does not succeed in identifying backdoored encoders created by IMPERATIVE. This is because the stealthy loss is designed to narrow the distribution gap between backdoored and normal data in both pixel and feature levels. This makes it challenging for DECREE to detect abnormal model internal representation changes when specific behaviors are activated.
@@
Beatrix. [48]  Beatrix [48] identifies poisoned samples by detecting abnormalities in the feature space. We used two pretraining datasets, CIFAR-10 and STL-10, and created backdoored encoders using BadEncoder and IMPERATIVE. By sampling 500 clean inputs and 500 poisoned samples, we applied Beatrix to differentiate them. The findings, as illustrated in Table 5, show that Beatrix effectively recognizes poisoned samples from BadEncoder with over 93% accuracy. However, Beatrix struggles to identify poisoned samples from IMPERATIVE, with a detection accuracy of below 50% on both CIFAR-10 and STL-10, which is like random guessing. We further analyze the reason why Beatrix fails to detect IMPERATIVE in Appendix C.1.1.
@@
Neural Cleanse. [83]  We then test the resilience of our method against a well-known reverse engineering defense, namely Neural Cleanse (NC) [83]. Specifically, it utilizes the anomaly index to measure the deviation of the reconstructed triggers by their sizes, labeling models with an anomaly index exceeding two as Trojan-infected. Since NC is specifically designed for classifiers and cannot be directly used on image encoders, NC is employed to identify backdoors in a compromised downstream classifier. Moreover, the testing dataset from a targeted downstream dataset serves as the clean dataset. Experimental outcomes, as illustrated in Table 3, indicate that NC is unable to detect the Trojan model created by our approach. While effective in identifying patch-based Trojans [44], this method presumes a uniform trigger pattern at the pixel level across different samples. Our method circumvents this by employing input-dependent triggers, meaning the pixel-level Trojan alterations vary across samples.
5.5 Generalization to Large-scale Dataset
We the generalization ability of our method to large-scale dataset by applying our attack to an image encoder pre-trained on ImageNet [60] released by Google [5]. We show the detailed experimental setup in Appendix D.6.4. Table 5 indicates that our method is highly effective on ImageNet. It achieves high ASR across different datasets, ranging from 96% to 100%, showing its capability to consistently deceive the target system. As shown in Table 3, the high SSIM values ranging from 0.989 to 0.998, high PSNR values ranging from 32.14 to 38.34, and low LPIPS values demonstrate that the perturbations made by IMPERATIVE are almost imperceptible, maintaining the visual quality of the images. Moreover, the BA is close to CA, indicating our attack maintains accuracy for the given downstream task despite the backdoor. Additionally, both our ASR and BA are much higher than those of the benchmark baseline ISSBA [41], proving ours has better performance.
@@ 6 Conclusions and Future Work
In this paper, we propose an imperceptible and effective backdoor attack against self-supervised models based on the optimized triggers that is disentangled to the augmented transformation in the self-supervised learning. Based on the evaluation across five different datasets and seven SSL algorithms, our attack is demonstrated to be both highly effective and stealthy. It also effectively bypasses existing backdoor defenses. For future work, it would be beneficial to expand the scope of research to include various other domains of machine learning, such as NLP and audio processing.
Appendix A More Attacking Cases of IMPERATIVE
A.1 Generalization to Other Representative SSL Frameworks
As elaborated on subsection 2.1, SSL frameworks for Contrastive Learning can be normally based on the following categories: negative examples, self-distillation, and clustering. To further evaluate the efficacy of IMPERATIVE, we test our method across the representative SSL algorithms in each category apart from the default SimCLR, i.e., negative examples (MoCo [22]), self-distillation (BYOL [17], SimSiam [6], and DINO [4]), and clustering (SwAV [3]). Table 7 illustrates the detailed parameters for different augmentation transforms used in the SSL algorithms, including Random resized crop, Random horizontal flip, Color jitter, Random gray scale, and Random polarization. The backbone feature extractor used here is ResNet18 [21]. Other experiment settings are the same as those in Appendix D.6.3.
Table 6 shows the evaluation results of applying IMPERATIVE to the encoders pre-trained utilizing various SSL algorithms on CIFAR-10. The SSL algorithms demonstrate high effectiveness in achieving successful backdoor attacks with ASR nearing or reaching perfection in many cases, with an average ASR of 99.54%. Additionally, the algorithms excel in maintaining the stealthiness of these attacks, evidenced by a high average SSIM of 0.95 and a high average PSNR value of 26.89, alongside a low LPIPS value of 0.011. These metrics across the SSL algorithms collectively indicate that while IMPERATIVE are highly effective at manipulating model behaviors, they do so in a way that remains largely undetectable through standard image quality assessments, which aligns with the conclusions drawn in section 5.
@@ A.1 Generalization to Other Representative SSL Frameworks
BYOL [17] STL10 39.31 98.33 0.95 26.94 0.0070
GTSRB 11.15 99.31 0.96 32.35 0.0020
SVHN 42.76 99.74 0.96 31.88 0.0024
@@
DINO [4] STL10 36.06 99.57 0.93 26.98 0.0141
GTSRB 11.49 99.83 0.95 27.11 0.0069
SVHN 37.66 99.99 0.93 25.18 0.0184
Average 50.32 99.54 0.95 26.89 0.0110
Table 6: Evaluations on other SSL algorithms, which are used to pre-train the encoders on CIFAR-10. IMPERATIVE keeps highly effective and stealthy across these algorithms.
Parameter SimCLR/CLIP MOCO SimSiam BYOL SwAV DINO Description
input_size 32/224 32 32 32 / / Size of the input image in pixels.
cj_prob 0.8 0.8 0.8 0.8 0.8 0.8 Probability that color jitter is applied.
cj_strength 1 0.5 1 1 0.5 0.5 Strength of the color jitter. cj_bright, cj_contrast, cj_sat, and cj_hue are multiplied by this value.
cj_bright 0.4 0.4 0.4 0.4 0.8 0.8 How much to jitter brightness.
cj_contrast 0.4 0.4 0.4 0.4 0.8 0.8 How much to jitter contrast.
cj_sat 0.4 0.4 0.4 0.2 0.8 0.4 How much to jitter saturation.
cj_hue 0.1 0.1 0.1 0.1 0.2 0.2 How much to jitter hue.
min_scale / 0.2 0.2 0.08 / / Minimum size of the randomized crop relative to the input_size.
random_gray_scale 0.2 0.2 0.2 0.2 0.2 0.2 Probability of conversion to grayscale.
vf_prob 0 0 0 0 0 0 Probability that vertical flip is applied.
hf_prob 0.5 0.5 0.5 0.5 0.5 0.5 Probability that horizontal flip is applied.
rr_prob 0 0 0 0 0 0 Probability that random rotation is applied.
crop_sizes / / / / [32,32] / Size of the input image in pixels for each crop category.
crop_counts / / / / [2,2] / Number of crops for each crop category.
crop_min_scales / / / / [0.14,0.14] / Min scales for each crop category.
crop_max_scales / / / / [32,32] / Max_scales for each crop category.
global_crop_size / / / / / 32 Crop size of the global views.
global_crop_scale / / / / / [0.4, 1.0] Tuple of min and max scales relative to global_crop_size.
local_crop_size / / / / / 96 Crop size of the local views.
local_crop_scale / / / / / [0.05, 0.4] Tuple of min and max scales relative to local_crop_size.
n_local_views / / / / / 6 Number of generated local views.
solarization_prob / / / / / 0.2 Probability to apply solarization on the second global view.
Table 7: Parameters of Augmentation Transforms in SSL Algorithms.
A.2 Generalization to Multi-modal Models
We utilize our IMPERATIVE to integrate a backdoor into the image encoder of the multi-modal model CLIP [58], which is comprised of both an image and a text encoder, pre-trained on 400 million internet-sourced (image, text) pairs. This backdoored image encoder is then employed to construct a multi-shot downstream classifier.
Due to the unavailability of CLIP’s original pre-training dataset, we utilize the training images from CIFAR-10, resized to 224×224×3, as our shadow dataset. We define $\alpha=0.1$ and $\beta=20$ in Algorithm 1, and $\mu=0.1$ in Algorithm 2 to normalize the loss terms across similar scales. Additionally, Algorithm 1 and Algorithm 2 are configured with a batch size of 16, and learning rates are set at $10^{-6}$ and 0.005, respectively, for a duration of 100 epochs. SVHN serves as the downstream dataset with the target label “one”. We employ the same reference inputs as described in [29]. All other experimental settings remain consistent with those outlined in Appendix D.6.3.
Table 8 presents the attack results of multi-shot classification using CLIP. IMPERATIVE achieves a similar high BA and ASR as BadEncoder [29] does, but it maintains much more stealthiness with high SSIM&PSNR and low LPIPIS values than BadEncoder which uses a white patch as the backdoor trigger.
Methods CA(%) BA(%) ASR(%) SSIM PSNR LPIPIS
@@ A.2 Generalization to Multi-modal Models
BadEncoder [29] 70.27 99.99 / / /
Table 8: Attack results on the multi-shot classification of CLIP. IMPERATIVE attains comparable high BA and ASR to those of BadEncoder [29], yet it offers significantly enhanced stealthiness.
Appendix B Ablation Study
As discussed in subsection 4.4, we utilize $\mathcal{L}_{disentangle}$ to enlarge the gap between the backdoored image distribution and the augmented image distribution. We use $\mathcal{L}_{alignment}$ to promote feature alignment between backdoored images and reference images. We use $\mathcal{L}_{stealthy}$ to make the backdoor trigger stealthy and natural. In Algorithm 1, we initialize a backdoor injector using a pre-trained one by Algorithm 2. We use STL10 as the dataset for pre-training encoders. In this section, we demonstrate the effects of each important component in subsection 4.4 by ablating them respectively.
As reported in Table 9, the ASRs of our method decrease across all target downstream datasets when ablating each component respectively. As reported in Table 10, the average SSIM and PSNR drop dramatically and LPIPS increases a lot across all datasets when ablating $\mathcal{L}_{stealthy}$. For example, in the ablation case of the target downstream dataset CIFAR10, the ASR decreases from 99.58% to 67.84%, 10.22%, 75.27% when ablating $\mathcal{L}_{disentangle}$, $\mathcal{L}_{alignment}$ and backdoor injector initialization respectively. This distinctly shows that the absence of key components significantly compromises the effectiveness of our backdoor attack methodology. Moreover, when ablating $\mathcal{L}_{stealthy}$, the SSIM drops from 0.99 to 0.1. The PSNR drops from 41.80 to 5.65. The LPIPS increases from 0.0002 to 0.55. This indicates that the backdoored images without the $\mathcal{L}_{stealthy}$ are more easily detectable and differ a lot from the original images.
@@ Appendix B Ablation Study
Table 9: ASR(%), BA(%) in ablation studies. Ablating the $\mathcal{L}_{\text{alignment}}$ means ablating the one in $\mathcal{L}_{\text{injector}}$ instead of $\mathcal{L}_{encoder}$, and ablating the pre-trained backdoor injector means randomly initializing it.
@@
Table 10: Results of ablating $\mathcal{L}_{stealthy}$. Images backdoored without the $\mathcal{L}_{stealthy}$ are more readily identifiable and exhibit significant deviations from the original images.
Appendix C More Evaluations
C.1 More Robustness Tests
C.1.1 Beatrix
@@
Further examination of Beatrix’s detection performance reveals that it determines whether a sample is poisoned based on the deviation in the feature matrix within the embedding space [48]. A sample is deemed poisoned if the deviation exceeds a specific threshold. Figure 4 displays the distribution of deviation values. Notably, poisoned samples by BadEncoder show significantly higher deviation than clean samples, enabling Beatrix to achieve high detection accuracy. Conversely, the deviation distributions between clean inputs and poisoned samples by IMPERATIVE overlap considerably, making them difficult to differentiate.
@@
Figure 4: Beatrix’s deviation distribution [48] for clean and backdoored data, where blue bars represent clean samples and orange bars indicate poisoned ones. Poisoned samples from BadEncoder display markedly higher deviations than clean samples, allowing Beatrix to detect them effectively. However, the deviation distributions for clean inputs and poisoned samples from IMPERATIVE significantly overlap, complicating their differentiation.
@@
Figure 5: Resilience to Grad-CAM [64]. The resemblance observed in these heatmaps indicates that IMPERATIVE is capable of resisting defenses that rely on Grad-CAM.
C.1.2 Grad-CAM

### v2

@@ 5.2 Effectiveness Evaluation
Effective Attack. As shown in Table 1, with different pre-trained and downstream datasets, our method achieves a high average ASR of 99.09% across various datasets. Additionally, Table 2 demonstrates that with the same pre-trained and downstream datasets, our approach also achieves nearly 100% ASRs. Our method outperforms all baseline methods in all scenarios, highlighting its robustness and superior effectiveness in executing successful backdoor attacks.
Accuracy Preservation. The downstream classifiers trained on the backdoored encoder maintain good accuracy on clean samples, as shown in Table 1. The average BA is 73.10% compared to the average CA of 72.96%, with the difference within 1%. This suggests that the backdoor introduced by our method does not compromise the classifier’s ability to label clean images correctly. This is because $\mathcal{L}_{\text{utility}}$ guarantees that the backdoored and clean image encoders yield similar feature vectors for clean inputs.
5.3 Stealthiness Evaluation
@@
Badencoder [27]/DRUPE [68] 0.8355 14.1110 0.07693 0.820 53.363
CTRL [34] 0.9025 32.4098 0.00034 0.865 71.138
WaNet [47] 0.7704 14.2372 0.07432 0.662 98.092
Ins-Kelvin [42] 0.4955 16.1925 0.14000 0.677 96.449
Ins-Xpro2 [42] 0.5981 17.9173 0.04434 0.817 35.084
POIENC [39] 0.1214 11.2787 0.15867 0.597 172.220
SSLBKD [56] 0.8737 16.2414 0.09640 0.891 118.320
BLTO [63] 0.8417 29.6756 0.00941 0.950 36.385
Ours 0.9633 35.8649 0.00896 0.969 16.320
Table 3: Stealthiness comparison to existing methods on CIFAR10. Our method remains stealthy. Detailed data are shown in Table A9.
Algorithmic Metrics. We first compare the average SSIM, PSNR, and LPIPS when the pre-trained dataset is CIFAR10 and downstream datasets are STL10, GTSRB, and SVHN injected with these backdoor triggers to compare the stealthiness of various backdoor attack methods. Table 3 indicates that our method exhibits strong stealthiness advantages with an average of 0.9633 SSIM, 35.8649 PSNR, 0.00896 LIPIS, 0.969 FSIM, and 16.320 FID indicating minimal structural changes to the images, hardly detectable noise, almost negligible perceptual difference between the original and perturbed images. Although CTRL achieves a better LPIPS, our method outperforms it in both SSIM and PSNR. Additionally, our average ASR is 99.09%, significantly higher than CTRL 29.32%(see Table 1), indicating that our method is more effective overall. More detailed data across various datasets, i.e., CIFAR10, STL10, GTSRB, SVHN, and ImageNet are shown in Table A9 and Table A10.
@@ 5.4 Robustness Evaluation
To assess the resilience of our method against current backdoor defenses, we deploy various SOTA backdoor defense strategies, i.e., DECREE [13], Beatrix [44], ASSET [48], Neural Cleanse (NC) [72], STRIP [14], Grad-CAM [58] for evaluation. Additionally, to further test the robustness of our method, we evaluate its endurance against the following commonly studied noises, i.e., JPEG compression [10, 11], Poisson noise [1, 79], and Salt&Pepper noise [35, 1]. We also design an adaptive defense method for INACTIVE. We show that INACTIVE cannot be defended by STRIP, NC, Grad-CAM, noises, and adaptive defense in subsection A5.1.
@@
DECREE. DECREE [13] identifies trojan attacks in pre-trained encoders by flagging an encoder as compromised if the reversed trigger’s $\mathcal{L}^{1}$ norm proportion falls below a 0.1 threshold. As shown in Table 4, the $\mathcal{P}\mathcal{L}^{1}$-Norm for each pre-trained and downstream dataset pair exceeds this threshold, so DECREE fails to detect backdoored encoders created by INACTIVE. This is because our invisible trigger breaks DECREE’s assumption of a visible patch trigger, and our stealthy loss further narrows the distribution gap between backdoored and normal data, masking internal model anomalies.
Pre-trained Dataset Downstream Dataset $\mathcal{P}\mathcal{L}^{1}$-Norm
CIFAR10 STL10 0.25
SVHN 0.39
GTSRB 0.15
STL10 CIFAR10 0.21
SVHN 0.34
GTSRB 0.20
Table 4: Evaluation results of DECREE [13]. A model is judged as backdoored if its $\mathcal{P}\mathcal{L}^{1}$-Norm <0.1.
@@
Beatrix. Beatrix [44] identifies poisoned samples by detecting abnormalities in the feature space. We use two pretraining datasets, CIFAR-10 and STL-10, and create backdoored encoders using BadEncoder and INACTIVE. By sampling 500 clean inputs and 500 poisoned samples, we applied Beatrix to differentiate them. We find (see Table 5) that Beatrix effectively recognizes poisoned samples from BadEncoder with over 93% accuracy. However, Beatrix struggles to identify poisoned samples from INACTIVE, with a detection accuracy of below 50% on both CIFAR-10 and STL-10, which is like random guessing. We further analyze the reasons for the defense failure in subsection A5.1.
Encoder Method TP FP FN TN Acc
CIFAR-10 BadEncoder 499 24 1 476 97.50%
Ours 0 24 500 476 47.60%
STL-10 BadEncoder 458 24 42 476 93.40%
Ours 5 24 495 476 48.10%
Table 5: Detection results by Beatrix [44]. It struggles to detect poisoned samples from ours.
ASSET. ASSET aims to distinguish between backdoored and clean samples by eliciting distinct behaviors in the model when processing these two data types, facilitating their separation [48]. We replicate their defensive techniques on our backdoored CIFAR-10 dataset. Specifically, we applied our synthesized trigger to CIFAR-10 (with a target label of 0) to create a poisoned version of CIFAR-10, maintaining a 100% poisoning rate as our default setting. The feature extractor used is the ResNet18 backbone, trained on this poisoned CIFAR-10 dataset.
The True Positive Rate (TPR) measures how effectively a backdoor detection method identifies backdoored samples, with a higher TPR (closer to 100%) indicating stronger filtering capability. The False Positive Rate (FPR) reflects the precision of this filtering: when TPR is sufficiently high, FPR shows the trade-off, highlighting the proportion of clean samples incorrectly flagged as backdoored. A lower FPR suggests fewer clean samples are mistakenly discarded, ensuring more clean data is retained for further use. Based on ASSET’s metrics, we calculated the TPR as 7.14% and the FPR as 1.8%, indicating that our poisoned data can largely evade ASSET’s detection.
5.5 Generalization to Large-scale Dataset
We assess the generalization of our method on a large-scale dataset by attacking an ImageNet-pre-trained encoder from Google [6]. We compare our method’s performance with ISSBA, which is also trained and tested on ImageNet in its paper. Experimental setups are detailed in subsection A6.5.
Experimental Results. Table 6 indicates that our method is highly effective on ImageNet, with an average 98.66% ASR across different datasets. Moreover, Table 7 indicates the high SSIM and PSNR values and low LPIPS values, demonstrating that the perturbations made by INACTIVE are almost imperceptible. Moreover, the average 83.91% BA is close to the average 83.59% CA, indicating our attack maintains accuracy for the given downstream task despite the backdoor. Additionally, both our ASR and BA are much higher than those of the baseline ISSBA [36], proving ours has better performance.
Downstream Dataset No Attack ISSBA [36] Ours
@@ 5.5 Generalization to Large-scale Dataset
STL10 95.68 92.58 9.97 93.48 100.00
GTSRB 80.32 66.29 5.10 82.84 96.00
SVHN 74.77 67.67 18.03 75.40 99.99
Average 83.59 75.51 11.03 83.91 98.66
Table 6: Comparative results (CA(%), BA(%), and ASR(%)) of ISSBA [36] and our attack on ImageNet. Ours constantly achieves the highest ASRs while maintaining accuracy on clean samples of the downstream classifiers.
[…]
ISSBA [36] 0.7329 31.3496 0.12424
Ours 0.9867 34.5733 0.01233
Table 7: Stealthiness comparison on ImageNet.
@@ 6 Conclusions and Future Work
In this paper, we propose an imperceptible and effective backdoor attack against self-supervised models based on the optimized triggers that are disentangled in the augmented transformation in the SSL. Based on the evaluation across five different datasets and six SSL algorithms, our attack is demonstrated to be both highly effective and stealthy. It also effectively bypasses existing backdoor defenses. For future work, it would be beneficial to expand the scope of research to include various other domains of machine learning, such as NLP and audio processing.
Supplementary Material
A1 Generalization to Other SSL Frameworks
SSL frameworks for Contrastive Learning can be normally based on the following categories: negative examples, self-distillation, and clustering. To further evaluate the efficacy of INACTIVE, we test our method across the representative SSL algorithms in each category apart from the default SimCLR, i.e., negative examples (MoCo [20]), self-distillation (BYOL [16], SimSiam [7]), and clustering (SwAV [4]). Table A5 illustrates the detailed parameters for different augmentation transforms used in the SSL algorithms. Other experiment settings are the same as subsection A6.5.
Experimental Results. Table A1 shows the evaluation results of applying INACTIVE to the encoders pre-trained utilizing various SSL algorithms on CIFAR-10. The SSL algorithms demonstrate high effectiveness in achieving successful backdoor attacks with ASR nearing or reaching perfection in many cases, with an average ASR of 99.62%. Additionally, the algorithms excel in maintaining the stealthiness of these attacks, evidenced by a high average SSIM of 0.95 and a high average PSNR value of 28.42, alongside a low LPIPS value of 0.0086. These metrics across the SSL algorithms collectively indicate that while INACTIVE is highly effective at manipulating model behaviors, they do so in a way that remains largely undetectable through standard image quality assessments.
@@
BYOL [16] STL10 79.49 99.98 0.96 31.44 0.0025
GTSRB 70.70 100.00 0.96 35.34 0.0011
SVHN 42.76 99.74 0.96 32.12 0.0024
@@
Average 64.72 99.62 0.95 28.42 0.0086
Table A1: Evaluations on other SSL algorithms, which are used to pre-train the encoders on CIFAR-10. INACTIVE keeps highly effective and stealthy across these algorithms.
[…]
@@
BadEncoder 70.27 99.99 0.28 9.98 0.6105
Table A2: Attack results on the multi-shot classification of CLIP. INACTIVE attains comparable high BA and ASR to those of BadEncoder [27], yet it offers significantly enhanced stealthiness.
A2 Generalization to Multi-modal Models
We utilize our INACTIVE to integrate a backdoor into the image encoder of the multi-modal model CLIP [53], which is comprised of both an image and a text encoder, pre-trained on 400 million internet-sourced (image, text) pairs. This backdoored image encoder is then employed to construct a multi-shot downstream classifier.
Experimental Results. Table A2 presents the attack results of multi-shot classification using CLIP. INACTIVE achieves a similar high BA and ASR as BadEncoder [27] does, but it maintains much more stealthiness with high SSIM&PSNR and low LPIPIS values than BadEncoder which uses a white patch as the backdoor trigger.
A3 Ablation Study
We utilize $\mathcal{L}_{\text{disentangle}}$ to enlarge the gap between the backdoored image distribution and the augmented image distribution. We use $\mathcal{L}_{\text{alignment}}$ to promote feature alignment between backdoored images and reference images. We use $\mathcal{L}_{\text{stealthy}}$ to make the backdoor trigger stealthy and natural. In Alg. 1, we initialize a backdoor injector using a pre-trained one by Alg. 2. We use STL10 as the dataset for pre-training encoders. In this section, we demonstrate the effects of each important component by ablating them respectively.
As reported in Table A3, the ASRs of our method decrease across all target downstream datasets when ablating each component respectively. As reported in Table A4, the average SSIM and PSNR drop dramatically and LPIPS increases a lot across all datasets when ablating $\mathcal{L}_{\text{stealthy}}$. For example, in the ablation case of the target downstream dataset CIFAR10, the ASR decreases from 99.58% to 67.84%, 10.22%, 75.27% when ablating $\mathcal{L}_{\text{disentangle}}$, $\mathcal{L}_{\text{alignment}}$ and backdoor injector initialization respectively. This distinctly shows that the absence of key components significantly compromises the effectiveness of our backdoor attack methodology. Moreover, when ablating $\mathcal{L}_{\text{stealthy}}$, the SSIM drops from 0.99 to 0.1. The PSNR drops from 41.80 to 5.65. The LPIPS increases from 0.0002 to 0.55 indicating that the backdoored images without the $\mathcal{L}_{\text{stealthy}}$ are more easily detectable and differ a lot from the original images.
@@
Table A3: ASR(%), BA(%) in ablation studies. Ablating the $\mathcal{L}_{\text{alignment}}$ means ablating the one in $\mathcal{L}_{\text{injector}}$ instead of $\mathcal{L}_{\text{encoder}}$, and ablating the pre-trained backdoor injector means randomly initializing it.
@@
Table A4: Results of ablating $\mathcal{L}_{\text{stealthy}}$. Images backdoored without the $\mathcal{L}_{\text{stealthy}}$ are more readily identifiable and exhibit significant deviations from the original images.
Parameter SimCLR/CLIP MOCO SimSiam BYOL SwAV Description
input_size 32/224 32 32 32 / Size of the input image in pixels.
cj_prob 0.8 0.8 0.8 0.8 0.8 Probability that color jitter is applied.
cj_strength 1 0.5 1 1 0.5 Strength of the color jitter. cj_bright, cj_contrast, cj_sat, and cj_hue are multiplied by this value.
cj_bright 0.4 0.4 0.4 0.4 0.8 How much to jitter brightness.
cj_contrast 0.4 0.4 0.4 0.4 0.8 How much to jitter contrast.
cj_sat 0.4 0.4 0.4 0.2 0.8 How much to jitter saturation.
cj_hue 0.1 0.1 0.1 0.1 0.2 How much to jitter hue.
min_scale / 0.2 0.2 0.08 / Minimum size of the randomized crop relative to the input_size.
random_gray_scale 0.2 0.2 0.2 0.2 0.2 Probability of conversion to grayscale.
vf_prob 0 0 0 0 0 Probability that vertical flip is applied.
hf_prob 0.5 0.5 0.5 0.5 0.5 Probability that horizontal flip is applied.
rr_prob 0 0 0 0 0 Probability that random rotation is applied.
crop_sizes / / / / [32,32] Size of the input image in pixels for each crop category.
crop_counts / / / / [2,2] Number of crops for each crop category.
crop_min_scales / / / / [0.14,0.14] Min scales for each crop category.
crop_max_scales / / / / [32,32] Max scales for each crop category.
Table A5: Parameters of Augmentation Transforms in SSL Algorithms.
A4 Proof
We aim to prove the following statement:
Theorem 3.1. Given a perfectly-trained encoder $\mathcal{F}_{\theta}$ based on the augmentations sampled from the predefined augmentation space $S_{A}$, it is impossible to inject a backdoor with a trigger function $\mathcal{T}\in S_{A}$.
Proof.
The encoder $\mathcal{F}_{\theta}$ is said to be perfectly-trained if it achieves maximal similarity on any pair of augmented versions of the same training sample $x$, for all augmentations in the augmentation space $S_{A}$. Formally, this means:
$s(\mathcal{F}_{\theta}(A_{1}(x)),\mathcal{F}_{\theta}(A_{2}(x)))=1,\quad\forall x\in\mathcal{X},\,\forall A_{1},A_{2}\in S_{A},$
where $s(\cdot,\cdot)$ denotes the similarity measure, which achieves its maximum value when the two inputs are identical in the feature space. As a result, for all $x\in\mathcal{X}$ and any $A_{1},A_{2}\in S_{A}$, the representations produced by $\mathcal{F}_{\theta}$ are identical:
$\mathcal{F}_{\theta}(A_{1}(x))=\mathcal{F}_{\theta}(A_{2}(x)).$
A backdoor injection introduces a trigger function $\mathcal{T}$ such that for any input $x$, the representation produced by $\mathcal{F}_{\theta}$ after applying the trigger, $\mathcal{F}_{\theta}(\mathcal{T}(x))$, maps to a target representation that is distinct from those of legitimate augmentations. In this case, we assume $\mathcal{T}\in S_{A}$, meaning the trigger function is an augmentation within the predefined augmentation space. Since $\mathcal{T}\in S_{A}$, by the definition of the perfectly-trained encoder, the following holds for all $x\in\mathcal{X}$ and any $A_{1},A_{2}\in S_{A}$:
$\mathcal{F}_{\theta}(A_{1}(x))=\mathcal{F}_{\theta}(\mathcal{T}(x))=\mathcal{F}_{\theta}(A_{2}(x)).$
This implies that the feature representation of the input $x$ after applying the trigger $\mathcal{T}$ is indistinguishable from those produced by any other augmentations in $S_{A}$. As such, the trigger $\mathcal{T}$ fails to produce a unique or distinct representation in the feature space. For a backdoor to be successfully injected, the trigger function $\mathcal{T}$ must produce a representation $\mathcal{F}_{\theta}(\mathcal{T}(x))$ that is distinct from those produced by other augmentations in $S_{A}$. However, this contradicts the property of the perfectly-trained encoder:
$\mathcal{F}_{\theta}(A_{1}(x))=\mathcal{F}_{\theta}(\mathcal{T}(x)),\quad\forall A_{1}\in S_{A}.$
Thus, $\mathcal{T}$ cannot create a distinct feature representation and, therefore, cannot function as a backdoor. We have shown that for a perfectly-trained encoder $\mathcal{F}_{\theta}$, trained on augmentations from $S_{A}$, any trigger function $\mathcal{T}\in S_{A}$ fails to produce a distinct representation necessary for a backdoor injection. Therefore, it is impossible to inject a backdoor with a trigger function $\mathcal{T}\in S_{A}$.
A5 More Evaluations
A5.1 More Robustness Tests
@@
Beatrix.  Further examination of Beatrix’s detection performance reveals that it determines whether a sample is poisoned based on the deviation in the feature matrix within the embedding space [44]. A sample is deemed poisoned if the deviation exceeds a specific threshold. Figure A1 displays the distribution of deviation values. Notably, poisoned samples by BadEncoder show significantly higher deviation than clean samples, enabling Beatrix to achieve high detection accuracy. Conversely, the deviation distributions between clean inputs and poisoned samples by INACTIVE overlap considerably, making them difficult to differentiate.
@@
Figure A1: Beatrix’s deviation distribution [44] for clean and backdoored data, where blue bars represent clean samples and orange bars indicate poisoned ones. Poisoned samples from BadEncoder display markedly higher deviations than clean samples, allowing Beatrix to detect them effectively. However, the deviation distributions for clean inputs and poisoned samples from INACTIVE significantly overlap, complicating their differentiation.
Neural Cleanse. We then test the resilience of our method against a well-known reverse engineering defense, namely Neural Cleanse (NC) [72]. Specifically, it utilizes the anomaly index to measure the deviation of the reconstructed triggers by their sizes, labeling models with an anomaly index exceeding two as Trojan-infected. Since NC is specifically designed for classifiers and cannot be directly used on image encoders, NC is employed to identify backdoors in a compromised downstream classifier. Moreover, the testing dataset from a targeted downstream dataset serves as the clean dataset. Experimental outcomes, as illustrated in Table A6, indicate that NC is unable to detect the Trojan model created by our approach. While effective in identifying patch-based Trojans [40], this method presumes a uniform trigger pattern at the pixel level across different samples. Our method circumvents this by employing input-dependent triggers, meaning the pixel-level Trojan alterations vary across samples.
Pre-trained Dataset Downstream Dataset Anomaly Index  [72]
CIFAR10 STL10 1.17
SVHN 1.42
GTSRB 1.61
STL10 CIFAR10 0.88
SVHN 1.23
GTSRB 1.43
Table A6: Evaluation results of neural cleanse [72]. A model is judged as backdoored if its anomaly index >2, so our attack cannot be detected by it.
@@
Figure A2: Resilience to Grad-CAM [58]. The resemblance observed in these heatmaps indicates that INACTIVE is capable of resisting defenses that rely on Grad-CAM.